OpenStudy (anonymous):

42.) the sum of a number and one-third of the number is equal to the sum of twice the number and ten. find the number??????????

OpenStudy (anonymous):

n+1/3n=2n+10 solve for n

OpenStudy (anonymous):

idk how to solve it w/the 1/3??? what do I do first on left side????

OpenStudy (anonymous):

You could start by finding the common denominator for all terms.

OpenStudy (anonymous):

6?

OpenStudy (anonymous):

6 or 20?

OpenStudy (anonymous):

Out of those two 6 would work the best, but 3 would probably your best choice.

OpenStudy (anonymous):

how would iine?????

OpenStudy (anonymous):

You multiply each of your terms by 3/3 except for your 1/3n term because it already has a denominator of 3.

OpenStudy (anonymous):

3/3 multiply by 2n and 10 too?????????

OpenStudy (anonymous):

\[3/3(n+1n/3=2n+10)\] \[3n/3 + n/3=6n/3+30/3\]

OpenStudy (anonymous):

Exclude the 1n/3 term in that distribution

OpenStudy (anonymous):

N=6n+10

OpenStudy (anonymous):

that's what I got, is it correct??

OpenStudy (anonymous):

Could you show me how you got that?

OpenStudy (anonymous):

3n/3+n/3=30/3 N=10

OpenStudy (anonymous):

Where is your 2n term?

OpenStudy (anonymous):

N=10+6n

OpenStudy (anonymous):

so the number is 10+6n ?????

OpenStudy (anonymous):

????????

OpenStudy (anonymous):

3n/3+n/3=6n/3+30/3 We could distibute a 3 to each of these terms so that they are no longer fractions. 3(3n/3+n/3=6n/3+30/3) 3n+n=6n+30

OpenStudy (anonymous):

and that equation might be easier to solve

OpenStudy (anonymous):

I got N=2n+30

OpenStudy (anonymous):

Can I ask where your getting the "N" in N=2n+30

OpenStudy (anonymous):

4n-4n=N

OpenStudy (anonymous):

Does N represent zero?

OpenStudy (anonymous):

oh yeah my bad/:

OpenStudy (anonymous):

If that is the case then your previous response "N=2n+30" is correct but still needs to be solved

OpenStudy (anonymous):

the number is 15 ????

OpenStudy (anonymous):

Indeed :)

OpenStudy (anonymous):

mind helping w/afew more pleaseeee???????????????

OpenStudy (anonymous):

How many?

OpenStudy (anonymous):

2-4 dif but similar on back

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

2.) separate 104 into two parts so that one part is 18 more than the other part. find each part???? @Tauist628

OpenStudy (anonymous):

104 split in two n+18 and n so 2n+18=104, solve for n

OpenStudy (anonymous):

I got N=43

OpenStudy (anonymous):

I got N=43 @Tauist628

OpenStudy (anonymous):

Correct :)

OpenStudy (anonymous):

but theres more than one unknown how do I find the 2nd unknown???? @Tauist628

OpenStudy (anonymous):

If you know what N is, then you should substitute it in for the equation of your 2nd unknown, N+18

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You're welcome.

OpenStudy (anonymous):

3 last qustions????

OpenStudy (anonymous):

Do they all have the same basic concept for solving them?

OpenStudy (anonymous):

yeah but 10-13 look harder

OpenStudy (anonymous):

Let's work them out together and such

OpenStudy (anonymous):

10.)the sm of the ages of Ruth and her mother is 75 years. The difference in their age is 29 years. how old is each?????? 11.) The sum of $180 was divided amoung 3 people so that the second person recievd $6 less than twice as much as te first, and the third received $7 more than the second. how much did each person receive????? 12.) The sum f $305 was divided among 3 people so that the secon received $4 ess than twice as much as the first, and the thirdrecieved $23 more than the second. HOW much did each person receive?????? 8.) Stanley worked 6 more than twice as many hou as Kathleen did.. How many hours did each work if tgether they worked 69 hours??????

OpenStudy (anonymous):

Well....

OpenStudy (anonymous):

10.) A system of equations should be set up. r+m=75 and m-r=29

OpenStudy (anonymous):

what number??????

OpenStudy (anonymous):

Solve for either variable Ex r=75-m or m=75-r for your first equation and m=29+r or r=m-29

OpenStudy (anonymous):

what number?????? @Tauista628

OpenStudy (anonymous):

@Tauist628

OpenStudy (anonymous):

What do you mean "what number"?

OpenStudy (anonymous):

what number was u even talking about??? ur last comment?

OpenStudy (anonymous):

OH, which question? I was doing question #10

OpenStudy (anonymous):

can u repeat whaever u said please?

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

10.) A system of equations should be set up. r+m=75 and m-r=29 Solve for either variable Ex r=75-m or m=75-r for your first equation and m=29+r or r=m-29

OpenStudy (anonymous):

how do u solve m=29+r?

OpenStudy (anonymous):

You plug into the original equation r+m=75 substitue "m" with (29+r) r+(29+r)=75

OpenStudy (anonymous):

solve for r

OpenStudy (anonymous):

I got r^2=75/29 ???

OpenStudy (anonymous):

You're not distributing the r, you're just adding r+(29+r)=r+r+29

OpenStudy (anonymous):

don't get it please help asap after this one I gtg

OpenStudy (anonymous):

ilike help on 11 and 12 if u can???? tho,

OpenStudy (anonymous):

r+r+29=75 2r+29=75 2r=46 r=23 Substitute this back in for your first equation m+r=75 m+23=75 m=52 so Ruth's age is 23 and her mother's age is 52

OpenStudy (anonymous):

Could those be done at another time?

OpenStudy (anonymous):

pleaseeee if u don't mind I could really use the help due tom. pleaseee

OpenStudy (anonymous):

I thought you said you had to go.

OpenStudy (anonymous):

yeah I got yelled at for being up at ten but have to finish my homework.. mind help????

OpenStudy (anonymous):

Okay could you retype #11 so I won't have to scroll as much

OpenStudy (anonymous):

11) the sumof $180 was divided among 3 people so that tcond perso received $6 less than twice as much as the first, and received $7 more than the secd. How much did each person receive?????

OpenStudy (anonymous):

PLease help ASAP

OpenStudy (anonymous):

I can't read the question very well. Which person is which?

OpenStudy (anonymous):

second person $6

OpenStudy (anonymous):

180=(2n-6)+(n-7)+n 180=4n-13 solve for n

OpenStudy (anonymous):

N=48.25

OpenStudy (anonymous):

now plug n in for your original equations 2n-6=2nd Person n-7=3rd Person n=1st Person

OpenStudy (anonymous):

not understanding

OpenStudy (anonymous):

2(48.25)-6=Money given to 2nd person 48.25-7=Money given to 3rd person 48.25=money given to first person

OpenStudy (anonymous):

ok thanks.. can u scroll up for 12 please????

OpenStudy (anonymous):

12.) The sum f $305 was divided among 3 people so that the secon received $4 ess than twice as much as the first, and the thirdrecieved $23 more than the second. HOW much did each person receive?????? Same concept 305=(2n-4)+(2n-4+23)+n 305=5n+15 290=5n solve for n then substitute back into your original equations 2(n)-4=Money given to 2nd person 2(n)-4+23=Money given to 3rd person (n)=Money given to 1st person

OpenStudy (anonymous):

lost????

OpenStudy (anonymous):

290=5n n=58 substitute back in 2(58)-4=Money to 2nd person 2(58)+19=Money to 3rd person 58=Money to 1st person Also, don't forget to put your labels

OpenStudy (anonymous):

thanks!!!!!!!!!!!

OpenStudy (anonymous):

Next question?

OpenStudy (anonymous):

that's it..thanks soooo much

OpenStudy (anonymous):

I would like to advise to make sure for #11 that each person is getting the correct corresponding value as I wasn't able to accurately tell which was which.

OpenStudy (anonymous):

ok...thanks!!!!!!

OpenStudy (anonymous):

If you ever need anymore help feel free to send a message and I'll help as much as I can. Bye